# Investigating VTubing as a Reconstruction of Streamer Self-Presentation: Identity, Performance, and Gender

### v1

@@ Abstract.
VTubers, or Virtual YouTubers, are live streamers who create streaming content using animated 2D or 3D virtual avatars. In recent years, there has been a significant increase in the number of VTuber creators and viewers across the globe. This practise has drawn research attention into topics such as viewers’ engagement behaviors and perceptions, however, as animated avatars offer more identity and performance flexibility than traditional live streaming where one uses their own body, little research has focused on how this flexibility influences how creators present themselves. This research thus seeks to fill this gap by presenting results from a qualitative study of 16 Chinese-speaking VTubers’ streaming practices. The data revealed that the virtual avatars that were used while live streaming afforded creators opportunities to present themselves using inflated presentations and resulted in inclusive interactions with viewers. The results also unveiled the inflated, and often sexualized, gender expressions of VTubers while they were situated in misogynistic environments. The socio-technical facets of VTubing were found to potentially reduce sexual harassment and sexism, whilst also raising self-objectification concerns.
@@
$c$s: Human-centered computing Empirical studies in collaborative and social computingccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
The adoption of a virtual avatar, however, can enable live streamers to create a new virtual identity that provides relative anonymity, affords alternate streaming performance experiences, and can enable for increased self-presentation opportunities including gender expression. The practise of using a 2D or 3D virtual avatar that is animated and voiced by a human actor (i.e., ‘Nakanohito’) is called Virtual YouTubing or VTubing. Since 2016, live streaming and video sharing platforms such as YouTube, Twitch, and Bilibili (a Chinese video-sharing and streaming platform) have seen an rapid increase in the number of Virtual YouTubers or VTubers that create and contribute content. Typical VTubers usually maintain a virtual identity to engage with viewers while live streaming, and often share singing or dancing videos using their virtual identities. Unlike traditional streamers, VTubers, notably in Asia, were known to be predominantly female. Prior work (Lu et al., 2021) from the viewer side also unveiled the prevalence of sexism and misogyny among audiences that migrated from the Otaku subculture (Ito et al., 2012) and virtual idol industry. How VTubers, notably female VTubers, present their identities via a virtual avatar, remains a significant yet unanswered question.
Before the emergence of VTubers, several researchers explored the opportunities that virtual avatars provided for self-presentation in avatar-based systems such as online games (e.g., World of Warcraft) and social Virtual Reality (VR; e.g., VRChat). This line of research highlighted how the proteus effect (Yee and Bailenson, 2007; Yee et al., 2009), i.e., a phenomenon where user behaviour could be influenced by the characteristics of an avatar, was a key aspect of self-presentation when using virtual avatars. While such systems have enabled users to experiment with new identities that are difficult to display in the physical world  (Stets and Burke, 2000; Huh and Williams, 2010; Bessière et al., 2007; Kim et al., 2012), user agency is controlled and limited by the choices made by the designers of avatar-mediated systems (Kolko, 1999; McDonough, 1999), including affordances of Character Creation Interfaces (CCIs) that support limited flexibility (Ducheneaut et al., 2009; McArthur et al., 2015) and rules of the virtual world that often dictates platform-specific self-presentation (Freeman et al., 2015; Freeman et al., 2016; Freeman and Maloney, 2021). VTubers can overcome the constraints of CCIs or virtual worlds by using design software to endow their avatars with complex human characteristic (e.g., rich facial expressions), and fine-tune their identities through live streaming for a long time beyond a one-shot customization. They can also actively engage and interact with viewers using their customized identities during or beyond live streams (e.g., Discord) without the constraints of a virtual world.
@@
The study data revealed the strategies that streamers use to customize and manage virtual identities that contain elements of their real self, and how such identities afford more exaggerated performances and more candid and inclusive social interactions between viewers and steamers, among others. The contributions of this research are thus (i) a nuanced understanding of how VTubers customize their virtual identity and engage with their audiences, as well as how viewer-streamer interactions are thus reshaped or extended, (ii) an exploration of the unique performance mechanisms that are used when virtual avatars are deployed while live streaming, which enhances our understanding of avatar-mediated self-presentation in social computing systems and (iii) the unique gender expressions of predominantly female VTubers when they are exposed to misogynist environments, which opens up opportunities for future CSCW research through feminist lens.
@@ 2.2. Self-Presentation in Avatar-Mediated Systems
The presentation of one’s self via a virtual avatar in a virtual world exemplifies the ‘proteus effect’ (Yee and Bailenson, 2007), where users are prone to altering their self-presentation based on inferences of their avatars’ expected behaviours. Yee et al.’s research found that such behavioural changes also translated to real-life face-to-face communication (Yee et al., 2009). Later research identified that users form attachments (Van Looy et al., 2012; Bopp et al., 2019; Birk et al., 2016) to their virtual avatars, especially when an avatar had a sense of personality, unique behaviours, intentions, and style. This effect suggests that a user might see their avatar as a second self in the virtual world that they would like to protect and may worry about (Schroeder, 2012), or become concerned if it was under attack or died (Wolfendale, 2007). More importantly, virtual avatars have been found to afford different social values and norms than those held in in real life. For example, Freeman et al.’s studies on in-game marriage in Audition demonstrated how the gender of an avatar in the game impacted users’ self-presentation during intimate experiences (Freeman et al., 2016; Freeman et al., 2015). They found that avatar-mediated relationships featured explicit performances of users’ self-development, sexuality, and ethnicity in the physical world. Based on data collected in Second Life, Yee et al. found that virtual avatars transferred gender impact on interpersonal distances in the offline world into virtual environments and that intimate virtual interactions still supported Equilibrium Theory (Argyle and Dean, 1965) in real life (Yee et al., 2007). In addition, virtual avatars enabled users to experiment with completely new identities (e.g., cross-gender play (Freeman et al., 2015; Freeman et al., 2016; Huh and Williams, 2010)), or acted as an reaffirmation of users’ existing identities (Yee et al., 2011; Ruberg and Shaw, 2017; Haimson, 2018). Ducheneaut et al. conducted an online survey of more than a hundred participants in 2009 and found that users of Maple Story, World of Warcraft, and Second Life were motivated to create digital bodies that were different from their own to realize an idealised self, stand out, or follow a trend (Ducheneaut et al., 2009).
@@
In these avatar-based systems, however, it is the designers that ultimately restrict the agency of a user (Kolko, 1999; McDonough, 1999). Although a user’s virtual avatar can be customized via a Character Creation Interface (CCI), these interfaces support limited flexibility (Ducheneaut et al., 2009; McArthur et al., 2015). Additionally, the performances of an avatar are also limited and affected by the design of the virtual world. For example, a study of social VR found self-presentation of users to be platform-specific (Freeman and Maloney, 2021), based on the social atmospheres afforded by the platform. Freeman et al.’s research on in-game marriage also observed user adaptations to the heterosexual view of marriage in Audition  (Freeman et al., 2015; Freeman et al., 2016). How avatar-mediated presentation would change if users, rather than platform designers, dictated identity customisation and performances remains an open research question. The present research aims to fill this gap by investigating the configuration of virtual avatars in live streaming settings, where creators have much more freedom.
@@ 2.3. Live Streaming as a New Medium of Self-Presentation
Since the late 2000s, live streaming has become increasingly popular across social media platforms and has been leveraged by users to build online communities and form social relationships  (Taylor, 2018; Consalvo, 2017). The affordances of live streaming have been intensively researched from the perspective of performance and interaction (Pellicone and Ahn, 2017; Kempe-Cook et al., 2019; Cai and Wohn, 2019; Wohn et al., 2018), content moderation (Cai and Wohn, 2019; Seering et al., 2017), privacy management (Li et al., 2018; Wu et al., 2022), and so on.Prior work has also found that the interactivity and sociality of live streaming has contributed greatly to its popularity (Hamilton et al., 2014; Lu et al., 2018b; Wulf et al., 2020).
@@ 3. Methodology
To gain deeper insights into VTubers’ identities, streaming practices, and self-presentation from third-person perspectives and creators’ subjective and conscious reflections, we conducted participant observation of 7 VTuber channels and semi-structured interviews with 9 VTubers (C1-7, I1-9; Table 1). As VTubing typically involves both video sharing and live streaming, all VTubers involved also regularly uploaded videos to their channels of themselves singing or dancing, highlights of their streams, etc. All VTubers were ethnically Chinese and regularly live streamed via YouTube Live, Twitch, or Bilibili. Bilibili is a Chinese video sharing platform similar to YouTube, that provides live streaming features and is famous for its Japanese ACG (anime, comics, and gaming) culture where VTubing originated.
VTuber ID Platform(s) Language Region
Avatar
Gender
VTuber
Gender
Affiliated with
a Company
Number of
Subscribers
Evaluation Method
@@ 3.1. Interviews
Purposive sampling (Robinson, 2014) was adopted during the recruitment of interviewees to ensure that participants came from diverse cultures, genders, and platforms. We reached out to approximately one hundred VTubers through e-mail or direct messaging, according to recommendations of streaming platforms and fan groups. All VTubers approached streamed regularly every week and had streamed for at least a month at the time they were contacted. A total of eight VTubers (6 female) agreed to be interviewed. The interviewees had live streamed as a VTuber for at least half a year and accumulated at least 3K subscribers at the time they were interviewed.
@@ 4.1.1. Construction Process
When a VTuber channel debuts, there will usually be an initial stream during which the VTuber introduces the identity of their avatar and demonstrates its upbringing, personality, hobbies, and so on, which partly indicates how the avatar will be performed by the ‘Nakanohito’. After the ‘debut’ stream, the virtual identity is further enriched and fine-tuned, both in its visual design and in how it is performed by the Nakanohito. For example, we observed that C1 and I5’s avatars put on a new Christmas-style outfit while streaming on Christmas Eve. I1 (female avatar & Nakanohito) reported that she once designed summer clothing by herself using Live2D because she “would like to change clothes with her fans in reality when the weather turns hot” (1(b)). She also claimed to have deliberately raised the pitch of her voice after streaming for a while because she felt it “sounds cuter” and “matches the 14-year-old identity”. I2 (female avatar & Nakanohito) said that her virtual identity, such as they way she dressed, her accessory choices, and her avatar’s personality was partly determined after discussions with her fans during live streams.
@@ 4.1.2. Blended Virtual and Real Identities
VTubers usually performed in a way that matched the expected visual and behavioral characteristics of their avatars, which differs from real-person streaming, which emphasises authenticity (Pellicone and Ahn, 2017; Tang et al., 2022). For example, C4’s (female avatar) virtual identity was to be an AI so she often acted as if she were an AI that spoke like a robot, switched between languages, and ridiculed the limits of humans. However, even when using such extreme personas, we found that such identities tended to blend with the real self of the VTuber. The design of the virtual avatar, along with its made-up virtual world and background, usually borrowed ideas from the real life or idealisations of the VTuber. For example, I1’s (female avatar & Nakanohito) avatar was designed to resemble a sparrow because it was quite commonplace in Hong Kong and she wanted to promote the culture of Hong Kong (1(a)). According to I2 (female avatar & Nakanohito), her identity used to be a herb fairy (1(c)) in a virtual world. She said she was not in good health at that time, so she claimed her identity was capable of curing all illnesses.
@@ 4.1.3. The Pseudonymity Afforded by Virtual Identities
(a) The identity of I1 was initially a 14-year-old and sparrow-looking girl. I1 hired an artist to design the avatar based on keywords such as 14-year-old, cute, talkative, etc. The avatar used accessories resembling sparrow wings to represent Hong Kong.
(b) I1 wanted to wear summer clothes with her fans in reality so she designed a summer outfit using Live2D.
(c) The identity of I2 was said to be a herb fairy that was capable of curing all illnesses. The avatar design, accessories, and personality was determined via discussion with her fans.
Figure 1. Visual representations of participant VTubers’ identity constructions. The three sketches were hand drawn based on their avatar appearances.
@@ 4.2.3. Inclusive Interaction
Despite VTubers being relatively outspoken and overacting, viewers have been found to be quite tolerant of such behaviors (Lu et al., 2021). I2 (female avatar & Nakanohito) stressed the tolerance of her viewers compared to real-person streaming experiences in game streams. She said that, “ The audiences become more tolerant … The viewers when I was a real-person streamer came to watch your stream because of the game instead of the streamer. They could blame you if you were bad, or simply made mistakes during the game, however, when streaming as a VTuber, my viewers would joke around inoffensively on the same occasion, and teach you how to play”. This also agrees with our observations. When C3 (female avatar) swore at her viewers during her stream, instead of taking offence, her fans joked around and asked her to swear at them (sometimes even harsher) for fun, knowing that C3 was acting.
@@ 5.1. Identity Co-Construction as a Form of Engagement
This research built upon prior knowledge about identity management in avatar-mediated systems and live streaming. Similar to other avatar-mediated systems such as social VR or virtual worlds (Freeman et al., 2015; Freeman et al., 2016; Huh and Williams, 2010; Ducheneaut et al., 2009), VTubing affords experimentation with virtual identities, which in turn affects the performances of Nakonohitos during live streams. In addition, as VTubers’ identities were often co-constructed between viewers and the Nakanohito, VTubers constantly fine-tuned their avatar design and performances based on viewer feedback and suggestions, or for the purpose of viewer engagement. Such co-construction constitutes a unique form of audience participation in live streams. Unlike previously studied viewer-streamer interaction such as chatting (Hamilton et al., 2014; Tang et al., 2016), polling (Cheung and Huang, 2011; Lessel et al., 2017), gifting (Hilvert-Bruce et al., 2018; Lu et al., 2018b), etc., VTubing directly enables viewers to shape the identities of streamers and how such identities should be performed. Such forms of interaction also complement the conceptualisation of co-performance found by Li et al. (Li et al., 2018) since viewer participation usually translates into performances of the avatar during live streams.
@@
Despite the exaggeration behaviors used,the VTubing community was generally inclusive and tolerant towards VTubers, which echos previous research on viewer perceptions of VTubers (Lu et al., 2021). Some interviewees (e.g., I2, I5) stressed the reduction in harassment and inclusiveness experienced during their VTubing practices. On the one hand, this could be attributed to the ‘sense of distance’ proposed by Lu et al. (Lu et al., 2021) that makes the Nakanohito invisible behind the scenes and reminds viewers that the inflated presentation is acting. On the other hand, the reduced non-verbal cues of VTubing may lead viewers to inflate and idealize their perception of a VTuber, as suggested by the hyper-personal model (Walther, 1996), which translates into a tolerant and inclusive environment.

### v2

@@ Abstract.
VTubers, or Virtual YouTubers, are live streamers who create streaming content using animated 2D or 3D virtual avatars. In recent years, there has been a significant increase in the number of VTuber creators and viewers across the globe. This practice has drawn research attention into topics such as viewers’ engagement behaviors and perceptions, however, as animated avatars offer more identity and performance flexibility than traditional live streaming where one uses their own body, little research has focused on how this flexibility influences how creators present themselves. This research thus seeks to fill this gap by presenting results from a qualitative study of 16 Chinese-speaking VTubers’ streaming practices. The data revealed that the virtual avatars that were used while live streaming afforded creators opportunities to present themselves using inflated presentations and resulted in inclusive interactions with viewers. The results also unveiled the inflated, and often sexualized, gender expressions of VTubers while they were situated in misogynistic environments. The socio-technical facets of VTubing were found to potentially reduce sexual harassment and sexism, whilst also raising self-objectification concerns.
@@
copyright: acmlicensedjournal: PACMHCIjournalyear: 2024journalvolume: 8journalnumber: CSCW1article: 80publicationmonth: 4price: 15.00doi: 10.1145/3637357ccs: Human-centered computing Empirical studies in collaborative and social computingccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
The adoption of a virtual avatar, however, can enable live streamers to create a new virtual identity that provides relative anonymity, affords alternate streaming performance experiences, and can enable for increased self-presentation opportunities including gender expression. The practise of using a 2D or 3D virtual avatar that is animated and voiced by a human actor (i.e., ‘Nakanohito’) is called Virtual YouTubing or VTubing. Since 2016, live streaming and video sharing platforms such as YouTube, Twitch, and Bilibili (a Chinese video-sharing and streaming platform) have seen an rapid increase in the number of Virtual YouTubers or VTubers that create and contribute content. Typical VTubers usually maintain a virtual identity to engage with viewers while live streaming, and often share singing or dancing videos using their virtual identities. Unlike traditional streamers, VTubers, notably in Asia, were known to be predominantly female. Prior work (Lu et al., 2021b) from the viewer side also unveiled the prevalence of sexism and misogyny among audiences that migrated from the Otaku subculture (Ito et al., 2012) and virtual idol industry. How VTubers, notably female VTubers, present their identities via a virtual avatar, remains a significant yet unanswered question.
Before the emergence of VTubers, several researchers explored the opportunities that virtual avatars provided for self-presentation in avatar-based systems such as online games (e.g., World of Warcraft) and social Virtual Reality (VR; e.g., VRChat). This line of research highlighted how the proteus effect (Yee and Bailenson, 2007; Yee et al., 2009), i.e., a phenomenon where user behaviour could be influenced by the characteristics of an avatar, was a key aspect of self-presentation when using virtual avatars. While such systems have enabled users to experiment with new identities that are difficult to display in the physical world  (Stets and Burke, 2000; Huh and Williams, 2010; Bessière et al., 2007; Kim et al., 2012), user agency is controlled and limited by the choices made by the designers of avatar-mediated systems (Kolko, 1999; McDonough, 1999), including affordances of Character Creation Interfaces (CCIs) that support limited flexibility (Ducheneaut et al., 2009; McArthur et al., 2015) and rules of the virtual world that often dictates platform-specific self-presentation (Freeman et al., 2015, 2016; Freeman and Maloney, 2021). VTubers can overcome the constraints of CCIs or virtual worlds by using design software to endow their avatars with complex human characteristic (e.g., rich facial expressions), and fine-tune their identities through live streaming for a long time beyond a one-shot customization. They can also actively engage and interact with viewers using their customized identities during or beyond live streams (e.g., Discord) without the constraints of a virtual world.
@@
The study data revealed the strategies that streamers use to customize and manage virtual identities that contain elements of their real self, and how such identities afford more exaggerated performances and more candid and inclusive social interactions between viewers and steamers, among others. The contributions of this research to HCI and CSCW are thus (i) a nuanced understanding of how VTubers customize their virtual identity and engage with their audiences, as well as how viewer-streamer interactions are thus reshaped or extended, (ii) an exploration of the unique performance mechanisms that are used when virtual avatars are deployed while live streaming, which enhances our understanding of avatar-mediated self-presentation in social computing systems and (iii) the unique gender expressions of predominantly female VTubers when they are exposed to misogynist environments, which opens up opportunities for future CSCW research.
@@ 2.2. Self-Presentation in Avatar-Mediated Systems
The presentation of one’s self via a virtual avatar in a virtual world exemplifies the ‘proteus effect’ (Yee and Bailenson, 2007), where users are prone to altering their self-presentation based on inferences of their avatars’ expected behaviours. Yee et al.’s research found that such behavioural changes also translated to real-life face-to-face communication (Yee et al., 2009). Later research identified that users form attachments (Van Looy et al., 2012; Bopp et al., 2019; Birk et al., 2016) to their virtual avatars, especially when an avatar had a sense of personality, unique behaviours, intentions, and style. This effect suggests that a user might see their avatar as a second self in the virtual world that they would like to protect and may worry about (Schroeder, 2012), or become concerned if it was under attack or died (Wolfendale, 2007). More importantly, virtual avatars have been found to afford different social values and norms than those held in in real life. For example, Freeman et al.’s studies on in-game marriage in Audition demonstrated how the gender of an avatar in the game impacted users’ self-presentation during intimate experiences (Freeman et al., 2016, 2015). They found that avatar-mediated relationships featured explicit performances of users’ self-development, sexuality, and ethnicity in the physical world. Based on data collected in Second Life, Yee et al. found that virtual avatars transferred gender impact on interpersonal distances in the offline world into virtual environments and that intimate virtual interactions still supported Equilibrium Theory (Argyle and Dean, 1965) in real life (Yee et al., 2007). In addition, virtual avatars enabled users to experiment with completely new identities (e.g., cross-gender play (Freeman et al., 2015, 2016; Huh and Williams, 2010)), or acted as an reaffirmation of users’ existing identities (Yee et al., 2011; Ruberg and Shaw, 2017; Haimson, 2018). Ducheneaut et al. conducted an online survey of more than a hundred participants in 2009 and found that users of Maple Story, World of Warcraft, and Second Life were motivated to create digital bodies that were different from their own to realize an idealised self, stand out, or follow a trend (Ducheneaut et al., 2009).
@@
In these avatar-based systems, however, it is the designers that ultimately restrict the agency of a user (Kolko, 1999; McDonough, 1999). Although a user’s virtual avatar can be customized via a Character Creation Interface (CCI), these interfaces support limited flexibility (Ducheneaut et al., 2009; McArthur et al., 2015). Additionally, the performances of an avatar are also limited and affected by the design of the virtual world. For example, a study of social VR found self-presentation of users to be platform-specific (Freeman and Maloney, 2021), based on the social atmospheres afforded by the platform. Freeman et al.’s research on in-game marriage also observed user adaptations to the heterosexual view of marriage in Audition  (Freeman et al., 2015, 2016). How avatar-mediated presentation would change if users, rather than platform designers, dictated identity customisation and performances remains an open research question. The present research aims to fill this gap by investigating the configuration of virtual avatars in live streaming settings, where creators have much more freedom.
@@ 2.3. Live Streaming as a New Medium of Self-Presentation
Since the late 2000s, live streaming has become increasingly popular across social media platforms and has been leveraged by users to build online communities and form social relationships  (Taylor, 2018; Consalvo, 2017). The affordances of live streaming have been intensively researched from the perspective of performance and interaction (Pellicone and Ahn, 2017; Kempe-Cook et al., 2019; Cai and Wohn, 2019; Wohn et al., 2018), cultural engagement (Lu et al., 2019b, 2021a), content moderation (Cai and Wohn, 2019; Seering et al., 2017), privacy management (Li et al., 2018; Wu et al., 2022), and so on. Prior work has also found that the interactivity and sociality of live streaming has contributed greatly to its popularity (Hamilton et al., 2014; Lu et al., 2018b; Wulf et al., 2020; Lu et al., 2019a).
@@ 3. Methodology
To gain deeper insights into VTubers’ identities, streaming practices, and self-presentation from third-person perspectives and creators’ subjective and conscious reflections, we conducted participant observation of 7 VTuber channels and semi-structured interviews with 9 VTubers (C1-7, I1-9; Table 1). The 7 observed VTuber (C1-7) and 9 interviewed VTubers (I1-9) are independent of each other. As VTubing typically involves both video sharing and live streaming, all VTubers involved also regularly uploaded videos to their channels of themselves singing or dancing, highlights of their streams, etc. All VTubers were ethnically Chinese and regularly live streamed via YouTube Live, Twitch, or Bilibili. Bilibili is a Chinese video sharing platform similar to YouTube, that provides live streaming features and is famous for its Japanese ACG (anime, comics, and gaming) culture where VTubing originated.
VTuber ID Platform(s) Language Region Avatar Gender VTuber Gender Affiliated with a Company Number of Subscribers Evaluation Method
@@ 3.1. Interviews
Purposive sampling (Robinson, 2014) was adopted during the recruitment of interviewees to ensure that participants came from diverse cultures, genders, and platforms. We reached out to approximately one hundred VTubers through e-mail or direct messaging, according to recommendations of streaming platforms and fan groups. All VTubers approached streamed regularly every week and had streamed for at least a month at the time they were contacted. A total of nine VTubers (6 female) agreed to be interviewed. The interviewees had live streamed as a VTuber for at least half a year and accumulated at least 3K subscribers at the time they were interviewed.
@@ 4.1.1. Construction Process
When a VTuber channel debuts, there will usually be an initial stream during which the VTuber introduces the identity of their avatar and demonstrates its upbringing, personality, hobbies, and so on, which partly indicates how the avatar will be performed by the ‘Nakanohito’. After the ‘debut’ stream, the virtual identity is further enriched and fine-tuned, both in its visual design and in how it is performed by the Nakanohito. For example, we observed that C1 and I5’s avatars put on a new Christmas-style outfit while streaming on Christmas Eve. I1 (female avatar & Nakanohito) reported that she once designed summer clothing by herself using Live2D because she “would like to change clothes with her fans in reality when the weather turns hot” (Figure 1). She also claimed to have deliberately raised the pitch of her voice after streaming for a while because she felt it “sounds cuter” and “matches the 14-year-old identity”. I2 (female avatar & Nakanohito) said that her virtual identity, such as they way she dressed, her accessory choices, and her avatar’s personality was partly determined after discussions with her fans during live streams.
@@ 4.1.2. Blended Virtual and Real Identities
VTubers usually performed in a way that matched the expected visual and behavioral characteristics of their avatars, which differs from real-person streaming, which emphasises authenticity (Pellicone and Ahn, 2017; Tang et al., 2022). For example, C4’s (female avatar) virtual identity was to be an AI so she often acted as if she were an AI that spoke like a robot, switched between languages, and ridiculed the limits of humans. However, even when using such extreme personas, we found that such identities tended to blend with the real self of the VTuber. The design of the virtual avatar, along with its made-up virtual world and background, usually borrowed ideas from the real life or idealisations of the VTuber. For example, I1’s (female avatar & Nakanohito) avatar was designed to resemble a sparrow because it was quite commonplace in Hong Kong and she wanted to promote the culture of Hong Kong (Figure 1). According to I2 (female avatar & Nakanohito), her identity used to be a herb fairy (Figure 1) in a virtual world. She said she was not in good health at that time, so she claimed her identity was capable of curing all illnesses.
@@ 4.1.3. The Pseudonymity Afforded by Virtual Identities
[…]
Figure 1. Visual representations of participant VTubers’ identity constructions. The three sketches were hand drawn by the research team based on observed avatar appearances.
@@ 4.2.3. Inclusive Interaction
Despite VTubers being relatively outspoken and overacting, viewers have been found to be quite tolerant of such behaviors (Lu et al., 2021b). I2 (female avatar & Nakanohito) stressed the tolerance of her viewers compared to real-person streaming experiences in game streams. She said that, “ The audiences become more tolerant … The viewers when I was a real-person streamer came to watch your stream because of the game instead of the streamer. They could blame you if you were bad, or simply made mistakes during the game, however, when streaming as a VTuber, my viewers would joke around inoffensively on the same occasion, and teach you how to play”. This also agrees with our observations. When C3 (female avatar) swore at her viewers during her stream, instead of taking offence, her fans joked around and asked her to swear at them (sometimes even harsher) for fun, knowing that C3 was acting.
@@ 5.1. Identity Co-Construction as a Form of Engagement
This research built upon prior knowledge about identity management in avatar-mediated systems and live streaming. Similar to other avatar-mediated systems such as social VR or virtual worlds (Freeman et al., 2015, 2016; Huh and Williams, 2010; Ducheneaut et al., 2009), VTubing affords experimentation with virtual identities, which in turn affects the performances of Nakonohitos during live streams. In addition, as VTubers’ identities were often co-constructed between viewers and the Nakanohito, VTubers constantly fine-tuned their avatar design and performances based on viewer feedback and suggestions, or for the purpose of viewer engagement. Such co-construction constitutes a unique form of audience participation in live streams. Unlike previously studied viewer-streamer interaction such as chatting (Hamilton et al., 2014; Tang et al., 2016), polling (Cheung and Huang, 2011; Lessel et al., 2017), gifting (Hilvert-Bruce et al., 2018; Lu et al., 2018b), etc., VTubing directly enables viewers to shape the identities of streamers and how such identities should be performed. Such forms of interaction also complement the conceptualisation of co-performance found by Li et al. (Li et al., 2018) since viewer participation usually translates into performances of the avatar during live streams.
@@
Despite the exaggeration behaviors used,the VTubing community was generally inclusive and tolerant towards VTubers, which echos previous research on viewer perceptions of VTubers (Lu et al., 2021b). Some interviewees (e.g., I2, I5) stressed the reduction in harassment and inclusiveness experienced during their VTubing practices. On the one hand, this could be attributed to the ‘sense of distance’ proposed by Lu et al. (Lu et al., 2021b) that makes the Nakanohito invisible behind the scenes and reminds viewers that the inflated presentation is acting. On the other hand, the reduced non-verbal cues of VTubing may lead viewers to inflate and idealize their perception of a VTuber, as suggested by the hyper-personal model (Walther, 1996), which translates into a tolerant and inclusive environment.